OpenStudy (anonymous):

Solve G = 5(h – 3j) for h.

OpenStudy (zzr0ck3r):

g = 5(h-3j) we want h on one side so distribute first to free up h g = 5h-15j now get down to the term with h...add 15j to both sides g+15j=5h now devide away the 5 (g+15j)/5 = h simplify g/5+3j = h

OpenStudy (anonymous):

h = the quantity G minus three j, divided by five h = the quantity G plus three j, divided by five h = the quantity G minus fifteen j, divided by five h = the quantity G plus fifteen j, divided by five mmm but these are the answers they give me

OpenStudy (zzr0ck3r):

the last one, they did not simplify to the last step that I did

OpenStudy (zzr0ck3r):

(g+15j)/5 = the quantity G plus fifteen j, divided by five

OpenStudy (anonymous):

ohh ok ! thanks !